AB INITIO DETERMINATION OF MOLECULAR PARAMETERS FOR ETHANE-LIKE MOLECULES

\underline{J.R. COOPER}, Department of Physics and Astronomy, University of Calgary, Calgary, Alberta T2N 1N4, Canada; L.-H. XU, Department of Physical Sciences, University of New Brunswick, Saint John, New Brunswick E2L 4L5, Canada; N. MOAZZEN-AHMADI, Department of Physics and Astronomy, University of Calgary, Calgary, Alberta T2N 1N4, Canada.

A new method has been developed for the ab initio determination of several vibration-torsion-rotational spectroscopic parameters for a symmetric top molecule with an internal rotor. In contrast to existing methodologies ^{a}^{,}^{b} which employ vibrational contact transformations either numerically or algebraically, the present model treats the molecule as vibrationally static but with a density distribution characteristic of the vibrational wavefunction. The second-order rotational constants A and B and torsional constant F, distortion parameters D_{J}, D_{K}, and D_{JK}, and torsional distortion parameters D_{m}, D_{Jm}, and D_{Km} have been determined for a series of ethane-like molecules from the results of ab initio calculations done at the CCSD(T) level. The technique has been applied to the molecules CH_{3}CH_{3}, CH_{3}CD_{3}, CD_{3}CD_{3}, and CH_{3}SiH_{3} with very promising results. Preliminary results for the potential constants F_{3J} and F_{3K} are also in excellent agreement with global fit values.

aY.-B. Duan, L. Wang, X.T. Wu, I. Mukhopadhyay, and K. Takagi, J. Chem. Phys. 111, 2385 (1999).

bT.J. Lukka and E. Kauppi, J. Chem. Phys. 103, 6586 (1995).